# Time Series Foundation Models: Benchmarking Challenges and Requirements

### v1

@@ 3.1 Challenge: Test Set Contamination Through Multi-Purpose Use of Datasets
Figure 1: Lineage of dataset collections (left) used for training and evaluating recent Time Series Foundation Models (right). Typically, a collection contains multiple different datasets. Lines indicate cases where at least one dataset of a collection was used for pre-training, train/test, or zero-shot evaluation.
@@
Our lineage analysis, visualized in Figure 1, reveals that each TSFM uses a different combination of pre-training, train/test, and zero-shot datasets.
While we did not find a TSFM that used the same set for different purposes, we still argue that this multi-purpose use of datasets is a problem. For example, suppose Model A uses a dataset for pre-training, and Model B uses the same dataset for evaluation. In that case, it becomes impossible to evaluate these models against each other on that dataset fairly.
A notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM). This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama).
Another interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations (indicated by green connections in Figure 1). However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime). Again, this means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection. Note that it is not easy to exactly determine which datasets a model has been trained on. It requires careful reading of the original paper and sometimes even analyzing published source code. Doing this for a dozen or more models in a benchmarking study requires substantial effort.
No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Table 2 and Table 3). Considering all investigated datasets, only 7% have never been used for pre-training or train/test fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
@@
Renaming datasets is another issue we encountered during our analysis. For example, the Monash dataset "Elecdemand" is a scaled subset ($1/1000$) of the "Australian Electricity Demand" dataset. Normalization practices could introduce unintended information leakage if these datasets are treated as separate. Similarly, the dataset "ElectricityLoadDiagrams20112014" appears under various names, such as "Electricity" in the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and "ECL" in the Informer (Zhou et al., 2021) collection. Dataset collections sometimes also contain the same variables but different observations. For instance, the car ride shares dataset is included in both the Monash and GluonTS collections. Although both repositories cite the same GitHub source [^4] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they represent different time series of distinct For-Hire Vehicles.
The empirical analysis using the Moirai TSFM by Aksu et al. (2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs. The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it. A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage. They then compared the performance of the models with different numbers of parameters and different forecasting horizons. The models trained on the training set with information leakage achieved an average MAPE between 7.6 and 59 percentage points lower than the same architecture trained on leakage-free training data (see Appendix B.2 for detailed results). These results vividly demonstrate the significant impact that test set contamination has on the performance of TSFMs. Larger model sizes especially seem prone to this, suggesting that bigger models tend to memorize rather than generalize. Another example of the information leakage effect is observable in a recently proposed TSFM benchmark. The benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b). Based on an analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets. The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%. Of course, the effect should be evaluated using comparisons on the same datasets, but the performance benefit through information leakage appears to align with findings from the Moirai leakage example. This highlights the difficulties for benchmarking between TSFMs.
3.2 Challenge: Memorization of Global Patterns
@@
Direct overlap between observations of the training and test sets is not the only cause for information leakage. A more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor & Narayanan, 2023). Typically, these indirect dependencies stem from confounding bias, that is, an external factor independently associated with observations in both datasets. In the context of time series forecasting, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains.
@@
A key strength of TSFMs is learning patterns and transferring a suitable pattern to an unseen time series. But when a TSFM is trained on data including such global events or crises, information may inadvertently carry over to forecasts on test set time series from the same period and affected by the same external factors. This introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation. As a result, the model gains access to information during training that is usually inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics. The time component and the actual information inside a time series plays a new role for globally trained models such as TSFMs.
Figure 2: Global pattern memorization of TSFM A when trained on time series 1 and zero-shot forecasting time series 2.
3.3 Challenge: Evaluation Dimensions
The typical evaluation in time series forecasting is performed across the time dimension to assess a model’s capabilities over time (Hyndman & Athanasopoulos, 2021), particularly its performance across different seasonal patterns (Figure 3, left side). Using time-series cross-validation, one can verify whether a model provides reliable forecasts, for example during winter and summer. It should be emphasized that a rolling evaluation does not necessarily require retraining the model for each rolling interval. One can just as well use a fixed training dataset and then roll inference over the various test sets (Hewamalage et al., 2023), making it a suitable evaluation strategy for TSFMs.
Figure 3: Evaluation strategies along the time or the domain dimension.
Interestingly, current TSFMs seem to evaluate only using a single time split per series (Figure 3, right side) without performing time series cross-validation, usually to reduce resource utilization or costs (Das et al., 2024; Gruver et al., 2024). Results are then often reported based on a single prediction per time series (Ansari et al., 2024; Das et al., 2024; Shi et al., 2024). They demonstrate foundation model capabilities as the models are evaluated against multiple time series from different domains (see Section 2). But we have not found any TSFM for which a combined time and domain cross-validation is explicitly conducted (Figure 4), even though this approach would easily increase the number of predictions and ultimately lead to a more meaningful evaluation.
Figure 4: Evaluation along time and domain dimension.
3.4 Challenge: Domain Representative Time Series
Foundation models in particular aim to be generalizable across different domains. Still, there is no systematic derivation of why domains such as energy, transportation, weather, or web trends in particular should be predicted. A particularly illustrative example is the ETT dataset (Zhou et al., 2021), which is used in many TSFMs as a zero-shot benchmark (see Section 3.1). However, this dataset is based on two years of measurements from just two transformers in China. The exact location of these transformers within China is not specified, and only very general information about their origin is provided. Moreover, while the dataset indeed captures some noteworthy characteristics, such as both short- and long-term seasonal patterns, it still consists of only two specific time-series use cases among an uncountable variety of possible temporal dynamics observed in real-world applications. It is therefore unclear to what extent results obtained on ETT can be considered generalizable, whether to transformers deployed in other regions such as Africa, America, Europe, or even to a third transformer within China. Accordingly, it remains questionable whether the dataset allows for any broader conclusions about the overall capability of models in a special domain (Hewamalage et al., 2023).
4 Current Benchmarking Approaches
One could argue that existing best practices for model training and testing are sufficient to mitigate risks such as test set contamination and global pattern memorization. But are these strategies still adequate in the age of TSFMs, or must they be adapted, combined, or restructured?
Existing strategies can be broadly categorized into benchmarking strategies and research best practices. As the dominant logic of time series benchmarking (Hyndman & Athanasopoulos, 2021), time-based train/test splits, including those in the Monash dataset collection (Godahewa et al., 2021) and TFB (Qiu et al., 2024), prevent leakage within a single TSFM. However, it is still unsuitable for fair benchmarking, as teams may choose different training and testing sets (see Section 2). Moreover, such predefined train/test splits typically do not use a global split point across all series in a collection, which would protect against information leakage through memorizing global patterns. For example, the Monash collection series uses varying start and end points, preventing the possibility of a global time-based split across all included time series (see Appendix 7).
In terms of foundation models, held-out datasets like GIFT-Eval (Aksu et al., 2024) focus on creating curated pre-training and test datasets without any overlap to prevent test set contamination. This approach enables a robust zero-shot forecasting evaluation, which is critical to TSFMs. However, the held-out dataset approach only applies to TSFMs developed in the future, as past models cannot be benchmarked reliably due to potential data contamination caused by overlap between their pre-training data and the held-out dataset. Furthermore, in case researchers want to include more data into pre-training than provided in a collection like GIFT-Eval, it requires careful checks for potential test set contamination in zero-shot forecasting. Lastly, global patterns like events or crises can still affect both the training and held-out test sets of such collection as long as no global time-based split is enforced.
Combining held-out datasets, global time-based train/test splits and retraining and benchmarking all TSFMs on a standard pre-training and test set would prevent information leakage and enable fair benchmarking. Yet, retraining all model architectures used in a benchmark is obviously very resource-intensive and prone to implementation errors. In addition, forcing all models to use the same data for pre-training mitigates the advantages of the scaling laws (Edwards et al., 2024).
Using anonymized real-world benchmarking test datasets (e.g., masking or scaling them) can also help to address information leakage concerns. Anonymizing real-world time series would help prevent the simple recall of original data points in case they have been included in a model’s pre-training set (Der et al., 2023). However, pre-processing techniques like scaling, feature engineering, and tokenization, which are usually implemented in a TSFM architecture, could potentially reverse such anonymization efforts, leading to the original or a very similar time series. Hence, the effectiveness of this approach is questionable and would require sophisticated anonymization techniques that preserve statistical properties while preventing direct recalling of data points.
Competition-based splits, like in the M-competitions (Makridakis et al., 2022), require forecasts on unpublished time-splits of known series. Such competitions start by publishing the training set while the test set remains private. Forecasts are registered during the competition, usually organized via Kaggle [^5] https://www.kaggle.com/competitions/m5-forecasting-accuracy/overview and metrics are calculated using the private test set. Although these approaches prevent test set contamination by registering forecasts based on time-based splits, they do not ensure a rigorous held-out dataset split for zero-shot forecasting. Furthermore, the hidden test set usually contains static, past data for evaluation throughout the competition. Besides enabling TSFMs to benefit from global pattern memorization, organizing such competitions requires substantial resources for administration, leading to long waiting periods between competitions. In a more innovative way, the recent M6 competition employed a strategy of using real-world live data from 100 financial assets and registering predictions into the real future (Makridakis et al., 2024). The main task was not forecasting but ranking the assets according to their future values. Still, it was a closed competition in a predefined time frame where long waiting period persists for results as with previous competitions which can be a hurdle in the fast developments of TSFMs.
In contrast, synthetic data for testing promises an unlimited amount of truly novel and so far unknown data. In fact, for pre-training, synthetic data has already been used successfully (Dooley et al., 2023). Yet, the external validity of solely using synthetic data for testing remains questionable, as synthetic data generation itself needs to be evaluated (Yuezhang et al., 2021).
Research best practices such as pre-registration for predictive modeling provide indirect mechanisms to mitigate risks of test set contamination due to multi-purpose use of datasets or global pattern memorization. Pre-registration for predictive modeling has been proposed as a rigorous research method originating from the experimental sciences (e.g., medicine, psychology) (Hofman et al., 2023). This method requires an ex-ante documentation and publication of the whole predictive modeling process before model training, including all planned data collection and preparation as well as modeling and evaluation activities. Model cards can complement this ex-post, by providing all information on data, training, and benchmarking to other researchers in one document (Mitchell et al., 2019). While these best research practices have the potential to prevent cherry picking of training and test sets and evaluation metrics and provide deeper insights into the models, they cannot mitigate the risks stemming from multi-purpose use of datasets and memorization of global patterns.
Lastly, the academic review process ensures proper data splits and detects any information leakage. While the peer review process has proven its ability to spot methodological issues, relying solely on this approach may not comprehensively prevent unintentional benchmark contamination (Li et al., 2024b). Furthermore, reviewers indeed cannot assess the validity of all models included in a benchmarking study. Hence, as witnessed in LLM evaluation research, academic review processes may oversee test set contaminations (see Section 2).
In sum, while pre-defined distinctions between training, test, and zero-shot sets improve evaluation integrity, they hinder comparability with prior models and mitigate scaling laws. Adhering to best practices may help identify test set contamination, but cannot entirely prevent it, especially in the presence of global pattern memorization. Given the rapid evolution of TSFMs, waiting long periods between competitions is increasingly impractical and risks stagnating meaningful progress, while anonymization of data cannot fully mitigate TSFM benchmarking challenges.

### v2

@@ 3.1 Direct Information Leakage: Train-Test Sample Overlap
Figure 2: Lineage of dataset collections (left) used for training and evaluating recent Time Series Foundation Models (right). Typically, a collection contains multiple different datasets. Lines indicate cases where at least one dataset of a collection was used for pre-training, train/test, or zero-shot evaluation. Smaller datasets (used n<10) are shown as "Other".
@@
Our lineage analysis, the details are described in Supplementary Information Section D and the results are visualized in Figure 2, reveals that there is no consensus between model developers which publicly available datasets should be used for what purpose. In fact, each TSFM team uses a different combination of pre-training, train/test, and zero-shot datasets.
Whereas we find no indication that any individual TSFM suffers from information leakage within its own training and evaluation, the widespread practice of re-purposing datasets introduces substantial challenges for benchmarking across TSFMs. Once a dataset has been used for pre-training or fine-tuning by any model, it must be regarded as contaminated and is no longer admissible for TSFM benchmarking. Our lineage analysis shows that this is not merely a theoretical concern, but a concrete and pervasive practical issue. No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Supplementary Information Section D). And only 6% of the analyzed datasets have never been used for pre-training or fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
In Supplementary Information Section B, we summarize literature showing that leakage between training and test sets can easily lead to over 50% better test scores. Several documented leakage cases suggest that the often unintended overlaps between train and test data are not always identified during the peer-review process Saravanan et al. (2024); Montet et al. (2025); Li et al. (2024b). This is not surprising as exactly determining on which datasets a model has been trained requires careful reading of the original paper, its appendix, and sometimes even analyzing published source code. Doing this for a dozen or more models when performing a benchmarking study requires substantial effort from the authors and reviewers as well. As more TSFMs are published, authors increasingly face the challenge of navigating complicated dataset relationships. For example, the authors of the Tirex model discovered that a dataset called "Solar" appears in parts in the Gift-Eval test data Aksu et al. (2024) as well as in the Chronos pre-training data, which is used by several models including Tirex, Toto, and Sundial Ansari et al. (2024a); Auer et al. (2025); Liu et al. (2025); Cohen et al. (2025). As a precaution, Tirex and Toto excluded the dataset, though Tirex explicitly notes the difficulty of verifying whether the data is actually identical. Through our lineage analysis, we found that in the Chronos pre-training data, the "Solar" dataset includes solar power data from all US states at two frequencies: 5T (5 minutes) and 1H (hourly), where the 5T data was sourced directly from the original data provider [^5] https://www.nlr.gov/grid/solar-power-data. Gift-Eval, however, sourced only the Alabama subset as a test set, aggregated to 10T, 1H, 1D (daily), and 1W (weekly), through the LSTNet dataset repository [^6] https://github.com/laiguokun/multivariate-time-series-data/tree/master/solar-energy. LSTNet also sourced the data from the original provider but, as noted, included only Alabama rather than all states. In addition, the Alabama solar subset appears in the Monash Repository [^7] https://forecastingdata.org/ at frequencies 10T and 1W. This modified version of the source data was used as pretraining data by the models TinyTimeMixers and Lag-Llama Ekambaram et al. (2024); Rasul et al. (2024). While TinyTimeMixers and Lag-Llama were published before Gift-Eval, the authors of Sundial validate against Gift-Eval and state they excluded any overlaps without explicitly listing which datasets they excluded. This example illustrates how dataset provenance can become convoluted through multiple transformations and repositories, making it difficult for researchers to identify overlaps.
@@
While in the previous example the dataset name was constantly "Solar", relying on dataset names to avoid train-test overlaps is not sufficient. The identification of suitable datasets is complicated by dataset remixing and renaming. For example, the dataset "Elecdemand" from the Monash Repository is a scaled subset ($1/1000$) of the “Australian Electricity Demand” dataset. Similarly, the dataset "ElectricityLoadDiagrams20112014" appears under various names, such as "Electricity" in the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and "ECL" in the Informer (Zhou et al., 2021) collection. Dataset collections sometimes also contain different observations of the the same variables. For instance, the car ride shares dataset is included with same frequency and from the same year 2015 in both the Monash and GluonTS collections. Although both repositories cite the same GitHub source [^8] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they contain different time series of distinct For-Hire Vehicles (Uber in GluonTS vs Lyft in Monash). Consequently it is necessary to always compare the underlying data itself, keeping in mind it could have been transformed in various ways (scaling, handling missing values, renaming, resampling, etc.).
3.2 Indirect Information Leakage: Temporal Overlap of Correlated Train and Test Series
@@
Direct overlap in data points between training and test sets is not the only cause for information leakage. In the context of time series data, a more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor and Narayanan, 2023). Often, these indirect dependencies stem from confounding bias, that is, an exogenous factor influencing observations from both datasets. For example, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains. However, not only global phenomena but also local factors can introduce such dependencies. Weather conditions at a specific location, for instance, directly influence solar energy production data, creating a causal relationship between meteorological measurements and power generation time series. Similarly, local traffic patterns may affect both air quality measurements and delivery time predictions within the same geographic area.
A key strength of TSFMs lies in their ability to learn subtle patterns in a flexible, data-driven manner and to apply these patterns to forecast future values of previously unseen time series. At large scale, however, TSFMs may inadvertently memorize patterns observed in one or more training time series and transfer them to series in the test set from the same time period. As long as training and test series are correctly ordered in time and do not overlap, this behavior does not constitute a serious issue. However, if training and test series overlap temporally and have a casual relationship (confounding or direct causation), such pattern transfer can result in information leakage. In the example illustrated in Figure 3, a highly flexible TSFM may learn the characteristic shape of the stock market decline caused by the global COVID-19 pandemic from the time series of the DAX index and reproduce it when tasked with forecasting the S&P 500 index over the same period.
Figure 3: Indirect Information Leakage - Temporal Overlap of Correlated Train Series A and Test Series B.
@@
This temporal overlap of correlated train-test samples introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation. As a result, the model gains access to information during training that would be inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics. Therefore the explicit time period of a pattern and its actually encoded information gain importance for globally trained models such as TSFMs.
This issue is not merely theoretical, but has been demonstrated empirically. Rodrigo and Ortiz (2024) conducted an experiment isolating the influence of temporal overlap information leakage. They build a setup with highly correlated time series, such as metro, bus, road, and train usage in Madrid for the year 2024. They trained two models in a global manner with all series. One model was trained with a common train-test split date for all series. For the other model only the metro time-series was cutted at the train-test split date, keeping the other series fully in the train set. After training, they only predicted the test period of the metro series, which was never part of the training dataset. The second model with full access to the other correlated series during training achieved a Mean Absolute Error (MAE) of 248,24 compared to 439,14 to the first model with a common train-test split date across all series. This corresponds to a performance improvement of approximately 43.47% attributable solely to this indirect form of leakage, confirming that access to correlated and temporally overlapping external information significantly distorts evaluation results even without direct access to the target series.
Figure 4: Zero-shot forecasting for S&P 500 stock index during the Covid shock in the year 2020. In blue is the ground truth, in green a pretrained model with data up to December 2019, in red a model pretrained until December 2020. In both cases S&P 500 is not part of the training data.
While the study by Rodrigo and Ortiz (2024) used time series where the complete training data is highly correlated and temporally overlapping with the complete test data, we replicate an TSFM training process in a more realistic scenario. We incorporate diverse data where only a small percentage of observations is affected by indirect leakage. To demonstrate the impact of temporal train-test leakage on model performance under these conditions, we conducted an experiment using a trained-from-scratch Time-Series Transformer (TST) (Zerveas et al., 2020). As we have seen in Section 3.1, any other pre-trained TSFM could be already affected by this form of information leakage. The encoder-only model shares multi-variate characteristics with recent models such as MOMENT or Chronos-2 (Goswami et al., 2024; Ansari et al., 2025) but is significantly smaller, enabling faster training saturation. We trained the model using 10 different random seeds for two different training sets [^9] Details on the general setup, used datasets, scaling, hyperparameter-tuning and results can be found in Supplementary Information Section C: one with a temporal overlap to the test set during the COVID-19 pandemic and one without such overlap. The first training set comprised the NN5 Daily and Tourism Monthly (Monash collection), and the American (Dow Jones), Brazilian (Bovespa), British (FTSE 100), Canadian (TSX), French (CAC 40), German (DAX) and Japanese (Nikkei 225) stock indices from 2015 to the end of 2019. The only difference in the second training set was that the stock index data spanned from 2016 to the end of 2020, comprising observations affected by the COVID-19 outbreak. No data points were shared directly between the training and test sets.
Table 2: Per-seed S&P 500 MAE breakdown for indirect leak from other countries stock indices vs no-leak experiments. The percentage difference shows how much better the model with leakage data performs than the non-leakage model. For each seed, the better MAE in displayed in bold.
Seed MAE Before Split MAE After Split
No Leak Indirect Leak % Diff No Leak Indirect Leak % Diff
410 230.53 227.72 1.22% 157.85 199.33 -26.28%
1425 219.68 178.09 18.93% 120.77 82.53 31.67%
1680 193.52 187.19 3.27% 80.74 81.22 -0.60%
1825 247.76 156.12 36.99% 110.47 70.36 36.31%
2287 231.19 206.28 10.78% 102.42 100.54 1.84%
3658 219.23 156.91 28.43% 117.25 84.62 27.84%
4013 224.75 193.16 14.06% 114.19 97.98 14.20%
4507 206.45 213.70 -3.51% 93.10 84.01 9.77%
8936 189.40 199.38 -5.27% 103.59 153.67 -48.35%
9675 206.30 182.99 11.30% 82.35 132.44 -60.82%
Mean 216.88 190.15 11.62% 108.27 108.67 -1.44%
Std Dev 17.15 21.88 21.10 38.83
We selected stock prices as our target variable as they are specifically hard to model appropriately in a long-term valid manner Goyal et al. (2024), suggesting that any performance gaps must come from external leaked information. For both models, we used the S&P 500 index as a zero-shot test set, which was not part of both datasets. Additionally, we chose a time-frame where a common cause had influence on many different time series and chose therefore the COVID-19 outbreak. In Figure 3, time series 1 shows the DAX index and time series 2 represents the S&P 500 index. Note how the COVID-19 outbreak is visible in both series.
Figure 4 displays the mean and standard deviation of the predictions for both models across the ten runs, and Table 2 shows the specific values for each seed. For almost all seeds, the model trained on data with temporal overlap to the test set predicted the initial market crash following the outbreak significantly better than the non-leakage model. In total, the temporal overlap model performed better eight out of ten times. Up to the first of May, the leakage model achieved an average MAE of about 190 compared to about 217 for the non-leakage model. For the remaining part, both models had a very similar average MAE, with the non-leakage model being slightly better (108.3 compared to 108.7 for the non-leakage model. When looking at the specific runs, shown in Table 2, the performance difference between the two models is up to about 37% in favor of the leakage model during the shock. Despite the temporal overlap representing only a small fraction of the overall training data, specifically seven time series out of 484 (or 1,600 out of 205,859 observations in total) (Table 6), this performance gap underscores the significant impact the small temporal overlap of correlated series can have on the predictive performance of a model.